Patients coming to a medical clinic have a mean weight of 207.6 pounds, with a standard deviation of 22.6 pounds. The distribution of weights is not assumed to be symmetric. Between what two weights does Chebyshev's Theorem guarantees that we will find at least 95% of the patients? Round your answers to the nearest tenth.
Answer #1

Solution :

Given that,  

mean = \mu = 207.6

standard deviation = \sigma = 22.6

Using standard normal table,

P(Z > z) =95 %

1 - P(Z < z) = 0.95

P(Z < z) = 1 - 0.95

P(Z < -1.65) = 0.05

z = -1.65

Using z-score formula,

x = z * \sigma + \mu

x = -1.65 * 22.6 + 207.6 = 170.3

The patients = 170.3